Solve system of linear equations by elimination
3x - 30 = y
7y - 6 = 3x

Answer by Boreal(15235)

3x-30=y
3x+6=7y, rewriting the second
subtract the second
3x-30=y
-3x-6=-7y
========
-36=-6y
y=6
That makes 3x-30=6
3x=36
x= 12
second one
7(6)-6=3(12)
42-6=36
(12,6)